OpenStudy (anonymous):

Art bought 4 boxes of pens. After Art gave his friend Ben 8 pens, he has a total of 112 pens. How many pens are in a box? Which of the following is the best equation for the problem?

OpenStudy (anonymous):

QUESTION 27 Art bought 4 boxes of pens. After Art gave his friend Ben 8 pens, he has a total of 112 pens. How many pens are in a box? Which of the following is the best equation for the problem? 4X - 8 = 112 4X + 8 = 112 X /4 - 8 = 112 112 - X/4 = 8

OpenStudy (anonymous):

laca i gave u the answer in the one where u listed all those problems

OpenStudy (anonymous):

send me the link

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

here i just copied the answers

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

q3 = B q4 = c q5 = c q6 = d q7 = c q8 = c q9 = a q10 = b q11 = b q12 = b q13 = a q14 = d q15 = a q16 = c q17 = c q18 = b q19 = c q20 = a q21 = a q22 = c q23 = d q24 = a q25 = c q26 = b q27 = a q28 = b q29 = a q30 = a q31 = c q32 = a q33 = d q34 = a q35 = a q36 = d q37 = b q38 = ? q39 = ? q40 = ? q41 = ? q42 = ? q43 = c q44 = d q45 = d

OpenStudy (anonymous):

38 - 42 did not have enough info or was just weird and confusing af